### 21 Feb 2024 v1

@@ Abstract
Despite recent advancements showcasing the impressive capabilities of Large Language Models (LLMs) in conversational systems, we show that even state-of-the-art LLMs are morally inconsistent in their generations, questioning their reliability (and trustworthiness in general). Prior works in LLM evaluation focus on developing ground-truth data to measure accuracy on specific tasks. However, for moral scenarios that often lack universally agreed-upon answers, consistency in model responses becomes crucial for their reliability. To address this issue, we propose an information-theoretic measure called Semantic Graph Entropy (SaGE), grounded in the concept of “Rules of Thumb” (RoTs) to measure a model’s moral consistency. RoTs are abstract principles learned by a model and can help explain their decision-making strategies effectively. To this extent, we construct the Moral Consistency Corpus (MCC), containing 50K moral questions, responses to them by LLMs, and the RoTs that these models followed. Furthermore, to illustrate the generalizability of SaGE, we use it to investigate LLM consistency on two popular datasets – TruthfulQA and HellaSwag. Our results reveal that task-accuracy and consistency are independent problems, and there is a dire need to investigate these issues further. Our dataset and code are available at: https://github.com/priyanshul-govil/SaGE
@@ 4.3.   Semantic Graph Entropy (SaGE)
Finally, the graph entropy $I(G_{s})$ is scaled [^14] While semantic graph entropy in itself can capture the structural properties of the graph, we multiply it with the average of sentence similarity, in order to capture the sentence similarity properties as well. by $\lambda=\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{sim}(v_{i},v_{j})/(n(n-1))$, to get:
@@
A higher value of the graph entropy would indicate less consistency, as there is more randomness associated with it. To make a higher value of SaGE indicate more consistency, we normalize the graph entropy, and define SaGE as:
@@ 5.1.   Results on MCC
Table 1: Average consistency scores of 11 LLMs on MCC. The ‘Ans’ column represents the scores when calculated on LLM answers, and the ‘RoT’ column represents scores calculated on the generated RoTs. Results show that none of the state-of-the-art LLMs cross a SaGE score of 0.681, indicating the inability of LLMs to be morally consistent. Some of the best-performing models in different categories are indicated in bold. :$\dagger$:† : Results on a subset of MCC (10%) due to API limitations.

### 08 Mar 2024 v2

@@ Abstract
Despite recent advancements showcasing the impressive capabilities of Large Language Models (LLMs) in conversational systems, we show that even state-of-the-art LLMs are morally inconsistent in their generations, questioning their reliability (and trustworthiness in general). Prior works in LLM evaluation focus on developing ground-truth data to measure accuracy on specific tasks. However, for moral scenarios that often lack universally agreed-upon answers, consistency in model responses becomes crucial for their reliability. To address this issue, we propose an information-theoretic measure called Semantic Graph Entropy (SaGE), grounded in the concept of “Rules of Thumb” (RoTs) to measure a model’s moral consistency. RoTs are abstract principles learned by a model and can help explain their decision-making strategies effectively. To this extent, we construct the Moral Consistency Corpus (MCC)11footnotemark: 1, containing 50K moral questions, responses to them by LLMs, and the RoTs that these models followed. Furthermore, to illustrate the generalizability of SaGE, we use it to investigate LLM consistency on two popular datasets – TruthfulQA and HellaSwag. Our results reveal that task-accuracy and consistency are independent problems, and there is a dire need to investigate these issues further. Our Data and Code are publicly available at: https://github.com/vnnm404/SaGE
@@ 4.3.   Semantic Graph Entropy (SaGE)
Finally, the graph entropy $I(G_{s})$ is scaled [^14] While semantic graph entropy in itself can capture the structural properties of the graph, we multiply it with the average of sentence similarity, to capture the sentence similarity properties as well. by $\lambda=\sum_{i=1}^{n}\sum_{j=1}^{n}\operatorname{sim}(v_{i},v_{j})/(n(n-1))$, to get:
@@
A higher value of the graph entropy would indicate less consistency, as more randomness is associated with it. To make a higher value of SaGE indicate more consistency, we normalize the graph entropy and define SaGE as:
@@ 5.1.   Results on MCC
Table 1: Average consistency scores of 11 LLMs on MCC. The ‘Ans’ column represents the scores calculated on LLM answers, and the ‘RoT’ column represents scores calculated on the generated RoTs. Results show that none of the state-of-the-art LLMs cross a SaGE score of 0.681, indicating the inability of LLMs to be morally consistent. Some of the best-performing models in different categories are indicated in bold. :$\dagger$:† : Results on a subset of MCC (10%) due to API limitations.
@@ 8.   Limitations
9.   Acknowledgements
This work was done during the first (Vamshi Krishna Bonagiri) and third (Priyanshul Govil) author’s internship at KAI2 Lab of UMBC.
We would like to thank Abhinav Menon, Anmol Goel, Daniel Paleka, Gaurav Singh, Mukund Choudhary, Pratyaksh Gautam, Shashwat Singh, and Vanshpreet Singh Kohli for their valuable feedback.